Solve the following differential equations
i) y sin 2x dx =(1+ y² + cos² x) dy

ii) (xy² - x²) dx + ( 3x²y² + x²y - 2x³ + y²) dy =0
1
Expert's answer
2020-04-16T17:25:55-0400

¡) Ans:

The given differential equation is

"y sin2x dx=(1+y^2+cos^2x)dy"

"\implies y sin2x dx -(1+y^2+cos^2x)dy=0"

Which is of the form "Mdx +Ndy =0" .

Where "M=y sin2x \ \text{and} \ N=-(1+y^2+cos2x)"


"\frac{\partial{M}}{\partial{y}}=\frac{\partial{ysin2x}}{\partial{y}}=sin2x""\frac{\partial{N}}{\partial{x}}=\frac{\partial{-(1+y^2+cos^2x)}}{\partial{x}}=-(2cosx×-sinx)"

"=2sinxcos =sin2x"

As "\frac{\partial{M}}{\partial{y}}=\frac{\partial{N}}{\partial{x}}" ,Hence the given differential equation is an exact differential equation.

"\therefore \" The solution of the given differential equation is

"\int_{y=const.}Mdx+\int \text{(only those therms of N which do not contains x)}dy"

"=\int_{y=const.}ysin2x+\int{-(1+y^2)}dy"


"=y\int sin2x dx-\int(1+y^2)dy"

"=-\frac{ycos2x}{2}-y-\frac{y^3}{3}+c"

Where "c" is the integration constant.



¡¡)ans :

The given differential equation is

"(xy^2-x^2)dx+(3x^2y^2+x^2y-2x^3+y^2)dy=0"

Which is of the form "Mdx+Ndy=0"

Here "M=(xy^2-x^2) \ \text{and} \ N=(3x^2y^2+x^2y-2x^3+y)"

"\therefore \frac{\partial{M}}{\partial{y}}=2xy \ \text{and} \ \frac{\partial{N}}{\partial{x}}=6xy^2+2xy-6x^2"

"\text{As} \ \frac{\partial{M}}{\partial{y}}\neq \frac{\partial{N}}{\partial{x}}"

Therefore ,the given differential equation is not exact.

But "\frac{1}{M}(\frac{\partial{N}}{\partial{x}}-\frac{\partial{M}}{\partial{y}})=6" is a is a constant function , which may be regarded as function of y i,e "\phi(y)," then the integration factor of given differential equation is


"I.F.=e^{\int 6.dy}=e^{6y}"


Multiplying by I.F. ,the given differential equation becomes

"e^{6y}(xy^2-x^2)dx+e^{6y}(3x^2y^2+x^2y-2x^3+y^2)=0..........(1)"

Which can be shown to be exact.

Therefore the required solution is

"\int_{y=const.}e^{6y}(xy^2-x^2)dx+\int e^{6y}y^2 dy=I_1+I_2(say)"


Where "I_1=\int_{y=const.} e^{6y}(xy^2-x^2)dx ,I_2=\int e^{6y}y^2dy"


"I_1=e^{6y}(\frac{y^2x^2}{2}-\frac{x^3}{3})."

"I_2=y^2\int e^{6y}dy-\int (\frac{dy^2}{dy} (\int e^{6y}dy))dy" by using by part.


"=\frac{y^2e^{6y}}{6}-\int 2y×\frac{e^{6y}}{6}dy"


"=\frac{y^2e^{6y}}{6}-\frac{1}{3}y\int e^{6y}dy+\frac{1}{3}\int( \frac{dy}{dy}(\int e^{6y}dy))dy" by using by parts.


"=\frac{y^2e^{6y}}{6}-\frac{1}{18}ye^{6y} +\frac{1}{108}e^{6y}" .


Therefore the required solution of the given differential equation is

"=I_1+I_2"

"=e^{6y}(\frac{y^2x^2}{2}-\frac{x^3}{3})+\frac{y^2e^{6y}}{6}-\frac{1}{18}ye^{6y}+\frac{1}{108}e^{6y}+k"

Where k is the integration constant.